OpenStudy (anonymous):

if you deposit $1,000 in an account that pays 6% annual interest compounded continuously, what will the balance be after 5 years A 1,932.06 B 1,349.86 C 29,803.91 D 1,232.06

OpenStudy (campbell_st):

A = p(1+r/100)^n so A = 1000(1+6/100)^5 A is the future value... I'll leave it to you to calculate

OpenStudy (anonymous):

i think this formula is applicable  considering that it's compounded continously. .

OpenStudy (anonymous):

what would be e

OpenStudy (anonymous):

do you have scientific calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

plug this in your calculator. .

OpenStudy (anonymous):